### v3

# Fine Tuning vs. Retrieval Augmented Generation for Less Popular Knowledge

Abstract.
Language Models (LMs) memorize a vast amount of factual knowledge, exhibiting strong performance across diverse tasks and domains. However, it has been observed that the performance diminishes when dealing with less-popular or low-frequency concepts and entities, for example in domain specific applications. The two prominent approaches to enhance the performance of LMs on low-frequent topics are: Retrieval Augmented Generation (RAG) and fine-tuning (FT) over synthetic data. This paper explores and evaluates the impact of RAG and FT on customizing LMs in handling low-frequency entities on question answering tasks. We conduct extensive experiments on twelve LMs of varying size and type and different fine tuning, data augmentation, and retrieval models. Our findings indicate that while FT boosts the performance across entities of varying popularity, RAG surpasses FT by a large margin particularly for least popular factual knowledge. Additionally, the success of both RAG and FT approaches is amplified by improving retrieval and data augmentation techniques. Fine tuning, while beneficial for small LMs, requires extensive resources. To address this issue, we propose the new Stimulus RAG approach that surpasses the effectiveness of fine tuning based approaches, thereby eliminating the need for the costly data augmentation and fine tuning step for enriching LMs with less popular factual knowledge.
Retrieval Augmented Generation, Fine Tuning, Data Augmentation
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Proceedings of the 2nd International ACM SIGIR Conference on Information Retrieval in the Asia Pacific 2024; December 9-12, 2024; Tokyo, Japanccs: Information systems Question answeringccs: Information systems Novelty in information retrievalccs: Information systems Language models
1. Introduction
Figure 1. Comparison of RAG and fine-tuning on StableLM2 performance in question answering over factual knowledge. RAG-based approaches significantly enhance the performance of the vanilla StableLM2, outperforming fine-tuning by a large margin. Our proposed SRAG approach outperforms all models, including the fine-tuning based approaches.
Language Models (LMs) exhibit outstanding capabilities in executing tasks that demand extensive memorization of factual data (Chowdhery et al., 2023). However, their memorization capabilities are constrained when dealing with less frequent entities (Mallen et al., 2023; Gerritse et al., 2022; Kandpal et al., 2023; Sun et al., 2023), and even the largest models may encounter the well-known ”hallucination” problem (Shuster et al., 2021) and temporal degradation (Kasai et al., 2022). Consequently, when LMs are intended for deployment in less resourced domains, customization becomes imperative to ensure optimal performance. A common example is within the industrial setup, where chatbots or Question Answering (QA) systems need to accurately answer users’ questions about a proprietary knowledge graph or intra-company terminology with limited textual description (Zhang et al., 2024a; Soudani et al., 2023).
Retrieval-Augmented Generation (RAG) and Fine-Tuning (FT) stand out as two prominent approaches for adapting LMs to specific domains (Soudani et al., 2023; Mosbach et al., 2023; de Luis Balaguer et al., 2024; Ovadia et al., 2023). RAG retrieves relevant information from a document corpus and enhances LM’s response generation through the implementation of in-context learning (ICL) (Zhang et al., 2024a; Cuconasu et al., 2024). Conversely, FT approach updates model weights to become adept at recalling specific information and enhance its memorization capabilities during inference (Asai et al., 2024). In the context of less popular knowledge, where limited data is available, data augmentation methods are utilized to generate synthetic training data, serving as an initial step towards FT (Soudani et al., 2024b, a).
In this paper, we aim to understand which approach is more appropriate for customizing LMs for less resourced domains. Specifically, we seek to answer this research question: (RQ1) How RAG compares to fine-tuning for question answering over less popular factual knowledge and which factors affect their performance? To address this question, we conduct a comprehensive comparison of RAG and fine-tuning methods for less popular knowledge, assuming that textual descriptions, albeit limited, are available for a specific domain and application. We therefore collect Wikipedia documents related to QA datasets over long-tail entities and apply two methods of knowledge injections: parametric knowledge injection using FT and non-parametric knowledge injection using RAG. For the FT approach, the LM is fine-tuned with synthetically generated QAs from these documents using data augmentation approaches (Alberti et al., 2019; Ushio et al., 2023). For the RAG approach, we use retrievers to rank the most relevant documents for a query. We investigate how the effectiveness of these methods is affected by the following aspects: (i) fine-tuning method; i.e., full FT vs. parameter efficient fine-tuning (PEFT), (ii) data augmentation method, (iii) LM type and size; i.e., decoder only vs. encoder-decoder models and varying size, ranging from 80M to 11B parameters, and (iv) retrieval model performance. Through exhaustive experimentation on twelve LMs and different setup of fine-tuning, data augmentation, and retrieval models, we arrive to the following conclusions:
  • Fine-tuning method: Comparing full FT with PEFT (i.e., QLoRA (Dettmers et al., 2023)) for LMs with less than 2 billion parameters, full FT is more effective in the downstream task performance. PEFT, however, preserves the reasoning ability of LMs needed for RAG and outperforms full FT when the fine-tuned models used in combination with RAG (Table 2).
  • Data augmentation method: Comparing prompt-based QA generation with QA generation by a state-of-the-art fine-tuned model (Ushio et al., 2023), the prompt-based QA generation demonstrates better performance for the factual QA task. This suggests that large LMs have a superior ability to generate high-quality QA pairs, although the fine-tuned model produce a greater volume of data (Table 2).
  • LM type and size: Comparing decoder-only with encoder-decoder LMs (Flan-T5 models of various sizes), decoder-only models outperform encoder-decoder models of similar size. Interestingly, larger LMs generally do not benefit from fine-tuning, while smaller do. Therefore, a small fine-tuned LM with RAG can preform on par or better than a large LM; e.g., StableLM2 (1.6B) vs. Llama3 (8B) (Table 3).
  • Retrieval model: Comparing retrievers with varying performance in the RAG system, we observe that as the popularity of factual knowledge increases, the performance of the retriever decreases (Figure 7). Moreover, the performance of the RAG system increases by using higher performance retriever (Figures 1 and 8).
  • Fine-tuning vs. RAG: Comparing these two knowledge injection methods, RAG substantially outperforms fine-tuning. with an ideal retriever, fine-tuned LMs with RAG models either outperform or perform on par with vanilla LMs with RAG in all but one case (Figure 1).
While fine-tuning improves accuracy in answering factual questions, both with and without RAG, it demands a considerable amount of effort and resources. This leads us to our second research question: (RQ2): Can we avoid the cost of fine-tuning by developing an advanced RAG approach that surpass the performance of fine-tuned LM with RAG? To answer this, we develop Stimulus RAG (SRAG), a new RAG approach that stimulates an LM to generate the correct response based on the provided hint in the prompt. The hint is extracted from the top retrieved documents by the retrieval model. Our results demonstrate that Stimuls RAG outperforms all other combinations of fine-tuning, both with and without retrieve-the-generate RAG.
To summarize, this paper makes the following contributions:
  • We study the effectiveness of fine-tuning and RAG approaches for question answering over less popular factual knowledge and compare the performance of these models across distinct setups: vanilla and fine-tuned models, both with and without RAG.
  • We perform extensive experiments to understand how FT and RAG models are affected by four different factors: data augmentation method, fine-tuning method, LM type and size, and retrieval model.
  • We propose a new RAG approach Stimulus RAG that outperforms all RAG and fine-tuning setups, thereby bypassing the need for expensive fine-tuning.
2. Related Work
Parametric and Non-parametric Knowledge. It is demonstrated that large pre-trained LMs memorize the significant amount of world knowledge in their parameters (parametric knowledge(Mallen et al., 2023). FT can update the parametric knowledge embedded in LMs and customize it for specific domain (Mosbach et al., 2023; Gerritse et al., 2022). One of the important principle for FT is data availability, which is limited specifically in specialized domains (de Luis Balaguer et al., 2024; Ovadia et al., 2023). Data augmentation (DA) addresses data scarcity problem by generating task- and domain-relevant samples from existing unlabeled texts. A common DA approach for the QA task is generating question-answer pairs through a four-step Pipeline, consisting of: passage selection, answer extraction, question generation, and consistency filtering (Alberti et al., 2019; Lewis et al., 2019, 2021; Ushio et al., 2023). Ushio et al. (2023) conducted an empirical study comparing three QA generation approaches: Pipeline, Multitask, and End-to-End (E2E) and showed the E2E approach outperforms others in downstream tasks.
Furthermore, a large body of work shows that augmenting LMs with nonparametric knowledge (i.e., retrieved text chunks) enables much smaller models to match the performance of larger models (Lewis et al., 2020). In this method, known as Retrieval Augmented Generation (RAG), an information retrieval system is utilized to find relevant documents and adds them to the input prompt to enhance response generation of LMs (Asai et al., 2023a, b).
As interest grows in refining pre-trained LMs for particular tasks, the comparison of FT and RAG strategies under equitable conditions is becoming increasingly important. Mosbach et al. (2023) explored the effectiveness of few-shot FT versus ICL for classification tasks in general domains. de Luis Balaguer et al. (2024) compared FT and RAG in answering agriculture and geography-specific questions. Ovadia et al. (2023) assessed the performance on multiple-choice questions in specialized areas like anatomy, astronomy, biology, and prehistory. In contrast to these studies, we directly address the integration of less popular factual knowledge into LMs, comparing various retrievers, data augmentation, and fine tuning methods.
Less Popular Knowledge. An entity’s popularity in LMs is gauged by its frequency in the model’s pre-training data (Godbole and Jia, 2023; Min et al., 2023), often assessed through the entity’s occurrences in a large corpus (Kandpal et al., 2023) via entity linking (van Hulst et al., 2020; De Cao et al., 2021). Due to the practical challenges of direct counting, e.g., annotation of large-scale collections with entities (Kamphuis et al., 2023), proxies are defined to approximate the popularity of factual knowledge. Sun et al. (2023) use traffic metrics and content density, while Maekawa et al. (2024) introduce the co-occurrence of the subject entity and relation predicate as a popularity proxy. Wikipedia pageviews are among the most prevalent methods for measuring the popularity of entities (Mallen et al., 2023; Sciavolino et al., 2021; Chen et al., 2021).
Figure 2. Overview of parametric and non-parametric knowledge injection for less popular factual knowledge. First, we prepare the corpus. Next, we generate knowledge in two formats: textual documents and synthetic QA pairs. Finally, we inject the knowledge into the prompt or LM parameters.
RAG Development. RAG introduces a new approach in AI, combining the strengths of both retrieval-based and generative models (Cuconasu et al., 2024). The concept of RAG was coined and popularized by Lewis et al. (2020), who introduced a model that combines a dense passage retriever with a sequence-to-sequence model, called Retrieve-then-Generate. This approach demonstrated substantial improvements in knowledge-intensive tasks. Several parameters affect a RAG system’s accuracy, including the performance of the retriever model (Cheng et al., 2023), the relevance of the passages included in the prompt context, their position, and their number (Cuconasu et al., 2024; Asai et al., 2024).
However, several works have argued that the Retrieve-then-Generate approach is not optimal for more complex tasks, and more advanced RAG systems are needed. Adaptive-RAG (Jeong et al., 2024) defines a classification-based RAG system to decide which RAG model should be used based on the question type. RAFT (Zhang et al., 2024a) trains a model to ignore documents that don’t help in answering the question, thereby adapting LMs to domain-specific RAG. Generate-then-Retrieve (GR) (Abbasiantaeb and Aliannejadi, 2024) argues that the Retrieve-then-Generate paradigm is insufficient when the answer must be obtained from multiple documents. They introduce the GR pipeline, which first generates multiple queries and then retrieves information for the generated queries. Mallen et al. (2023) found that for popular knowledge, using RAG can hurt performance, so they defined an adaptive retrieval system to use retrieval only where it is beneficial.
We discuss that increasing the number of documents is not helpful and may introduce more noise into the LM’s input. To address this problem, we propose a new stimulus RAG system that highlights parts of the input text most likely to contain the correct answer. This approach aids the LM in accurately identifying and extracting relevant information. Highlighting has been used in the literature of information retrieval for various purposes. Askari et al. (2023b) aim to generate synthetic documents for queries, highlighting keywords to create high-quality documents. Cho et al. (2020) propose generating sub-sentence summary highlights to overlay on source documents, enabling users to quickly navigate through content. Li et al. (2023) introduce a new prompting framework to provide black-box LMs with fine-grained, instance-specific guidance toward desired outputs. Our work differs from these in that we highlight sentences using a simple yet effective reranker model, which directly improves RAG’s performance.
3. Methodology
In this section, we introduce our evaluation framework (Figure 2), which is designed to assess the effectiveness of two knowledge injection methods: the parametric method using FT and the non-parametric method using RAG.
3.1. Task Definition
This study specifically focuses on factual knowledge (Adams, 2015) of entities, defined as information about particular attributes and characteristics of target entities, among various types of world knowledge (Mallen et al., 2023). We chose this because the amount of knowledge memorized by LMs can be approximated by its accuracy in answering simple factual questions, such as ”In what city was Lisa Miller born?” (Sun et al., 2023).
Factual knowledge is defined as a triplet of (subject, relationship, object) (Mallen et al., 2023). In this context, the question involves the subject and the relationship, while the answer corresponds to the object. By using this template questions, we ensure that LMs understand the question and derive the answer from their embedded knowledge. We select Wikipedia-based question-answering datasets focused on factual knowledge. This enable us to measure the popularity of entities based on Wikipedia pageviews and also obtain the corresponding evidence document for each entity from Wikipedia. For each dataset, we select Wikipedia pages whose corresponding entities appear in the test dataset. This setup mirrors real-world industry practices, where entities and their textual descriptions relate to companies’ specific internal concepts.
3.2. Knowledge Injection with Fine-tuning
LMs are primarily pre-trained on general domains. To customize LMs for specific knowledge or a particular domain, fine-tuning (FT), also known as parametric knowledge injection, is commonly used. However, FT requires a substantial amount of training samples, which are often unavailable for specific applications, such as those within a company. Data augmentation (DA) offers a solution to this training data shortage. To achieve parametric knowledge injection, we fine-tune LMs using synthetically generated question-answer (QA) pairs. Formally, given a set of documents $D=\{d_{1},\ldots,d_{n}\}$, where $n$ is the number of documents, a QA pair generator $\mathcal{G}_{qa}$ is tasked with generating as many QA pairs as possible:
$\mathcal{G}_{qa}(d_{i})=\{(q_{i}^{1},a_{i}^{2}),\ldots,(q_{i}^{m},a_{i}^{m})\},$
where $q_{i}$ and $a_{i}$ denote a question and answer generated for the document $d_{i}$, $m$ is the total number of generated QAs for the document, and the generated set of question-answer pairs is denoted as $Q_{i}$. The QA generation method $\mathcal{G}_{qa}$ is applied to all documents $D$, and the generated set $Q=\bigcup_{i=1}^{n}Q_{i}$ is then used for fine tuning an LLM, which consists of a set of learnable parameters to predict the probabilities of future or masked tokens.
Figure 3. Input prompt for prompt-based QA pair generation. We define a CoT prompt to outline the generation steps.
We employ two QA generation methods. The first one is end-to-end (E2E) QA generation, where a fine-tuned sequence-to-sequence model generates QA pairs from $d_{i}$. We utilize the E2E approach by  Ushio et al. (2023), which employs a trained T5-large for paragraph-level QA generation and has been shown to be more robust and effective than the established pipeline approach. The term ”E2E” is used because the QA generation process is not divided into two sequential components, i.e., answer extraction and question generation. Instead, a QA pair is generated in one go. To train the E2E model, the training question-answer pairs $Q^{train}_{i}$ corresponding to the training document $d^{train}_{i}$ are converted into a flattened sentence $y_{i}$ the following transformation:
$\mathcal{T}(q,a)=\text{``{question:} }\{q\},\text{ {answer:} }\{a\}\text{''}$ ”
$y_{i}=``\{\mathcal{T}(q^{1}_{i},a^{1}_{i})\}\mid\{\mathcal{T}(q^{2}_{i},a^{2}_{i})\}\mid\ldots\text{''},$ ” ,
where each pair is textualized with the function $\mathcal{T}(q,a)$, and the textualized QA pairs are concatenated using the separator —. The E2E QA generation function $\mathcal{G}_{qa}$ is then obtained by maximizing the conditional log-likelihood:
$\arg\max_{y}P_{qa}(y\mid d).$
The second QA generation method is the prompt approach, in which the QA pair generator $\mathcal{G}_{qa}$ is an instruction-tuned LM capable of reasoning over the input prompt $I$. In this approach, the QA pairs are generated as follows: $\mathcal{G}_{qa}(d_{i})=LM(d_{i},I).$ We utilize Zephyr (Tunstall et al., 2023) with Chain of Thought (CoT) (Wei et al., 2022) prompting for QA generation, as demonstrated in Figure 3.
3.3. Knowledge Injection with RAG
The non-parametric knowledge injection is performed using RAG, which consists of two components: the Retriever and the Generator (Cuconasu et al., 2024; Asai et al., 2024).
Retriever. The first key component in a RAG system is a retriever $R$, which builds an index for a document corpus $D$. During inference, given an input sequence $q$, the retriever identifies and ranks relevant documents $D_{q}=R(D,q)$. In our retrieval process, we employ both sparse and dense retrievers. We utilize BM25 (Robertson and Walker, 1994) as a sparse retriever due to its popularity and effectiveness. For dense retrievers, we employ DPR (Karpukhin et al., 2020) and Contriever (Izacard et al., 2022) methods. Both models convert textual data into vector representations using a transformer network. The similarity between the query $q$ and document $d$ is defined as $S(q,d)=\vec{q}\cdot\vec{d}$, which computes the dot product between embedding vectors $\vec{q}$ and $\vec{d}$. DPR employs two independent BERT models, trained discriminatively using query-documents pairs with negative samples from BM25. Contriever, on the other hand, is trained using a shared BERT model for query and document encoding, optimized using a contrastive loss. We also employ a two-step retrieval pipeline, which includes first-stage retrieval using BM25 and reranking using DPR (Askari et al., 2023a; Lin et al., 2021; Abbasiantaeb and Aliannejadi, 2024).
Generator. The second step involves a generator component responsible for synthesizing an answer, typically implemented via LMs. Generative LMs operate by predicting the probability distribution of the next token, given the previous tokens. In RAG, the generative LM takes a query $q$ and top-$K$ ranked documents from $D_{q}$, denoted as $D_{q}^{K}=[d_{1},...,d_{K}]$, and generates a response by sequentially predicting the next token. Our RAG prompt prepends the documents before the query, following (Mallen et al., 2023; Cuconasu et al., 2024).
In this paper, we define and assess four distinct configurations of injecting knowledge with fine tuning and RAG: (1) -FT-RAG: the vanilla LM without retrieved documents, (2) -FT+RAG: the vanilla LM with retrieved documents, (3) +FT-RAG: the fine-tuned LM without retrieved documents, (4) +FT+RAG: the fine-tuned LM with retrieved documents.
Figure 4. Our proposed Stimulus RAG method. The Hint Extractor identifies the most relevant sentence from top-K documents ranked by the retriever. This sentence is then added to the beginning of the input prompt.
3.4. Stimulus RAG
While the generic retrieve-then-generate framework of RAG is effective in answering factual knowledge (Mallen et al., 2023), it sometimes struggles to accurately respond factual questions that are not memorized in LMs parameters, even when the ground truth document is included in the prompt. We hypothesize that by employing a more advanced RAG approach we can achieve or even surpass the benefits of the combined RAG and FT (+FT+RAG) setup. This would allow us to bypass the computationally expensive process of knowledge injection with FT, which involves resource-intensive processes of QA generation and FT.
We introduce Stimulus RAG, a RAG approach that guides LMs to generate responses using a hint provided in the RAG prompt. Stimulus RAG comprises three steps: Retriever, Hint extractor, and Generator. The Retriever and Generator are as defined in Section 3.3. The Hint extractor provide hint text that guides LM to generate accurate responses and works as follows: The top-$K$ ranked documents $D_{q}^{K}=[d_{1},...,d_{K}]$ ($d_{K}$ denoting the retrieved document at rank $K$) from the retriever step are split into sentences, denoted as $S_{q}=\{s_{j}\}_{j=0}^{N}$, where $N$ is the total number of sentences. These sentences are then ranked using a retrieval model $R^{\prime}(S_{q},q)$. For the sake of simplicity, we use the same retrieval model as that used in the retriever step, hence $R=R^{\prime}$. The top-ranked sentence $s_{q}^{1}$ is identified as the most relevant sentence for the given question and serves as a hint. The sentence $s_{q}^{1}$ can be directly used in the prompt, which we refer to as the SRAG(S) approach. Alternatively, the document containing the sentence $s_{q}^{1}$ can be considered as a hint, which we refer to as the SRAG(D) approach. The extracted hint is placed at the top of the input prompt provided to the Generator component. This implies that the hint text a repetition of presumably the most relevant sentence/document of in the prompt. It has been shown that the beginning of the input prompt receives more attention from LMs (Liu et al., 2023). Our method is illustrated in Figure 4.
Table 1. Statistics of the factual knowledge-based datasets.
Dataset # QA # Rel. Type Question form
PopQA 14K 16 Template
WiTQA 14K 32 Model-assisted
EQ 17.3K 24 Template
Figure 5. Distribution of sample counts across popularity buckets, defined by $log_{10}(\text{pageviews})$ for PopQA and WitQA and $log_{2}(\text{pageviews})$ for EQ.
4. Experimental Setup
Table 2. Accuracy of vanilla and fine-tuned LMs, both with and without RAG. The RAG results are based on ideal retrieval. Statistically significant differences in the PEFT-Prompt rows are compared with other rows. Superscripts (a), (b), (c), and (d) indicate statistically significant differences (better or worse) compared to vanilla LM, PEFT-E2E, Full-E2E, and Full-Prompt, respectively, determined by the Wilcoxon test (p-value ¡ $0.01$).
PopQA EQ
FT QA +FT-RAG +FT+RAG +FT-RAG +FT+RAG
FlanT5-small 2.69 47.46 2.84 27.36
PEFT E2E 6.06 48.40 7.86 33.76
PEFT Prompt 7.01(a,c,d) 61.39(a,b,c,d) 9.39(a,b,c,d) 42.68(a,b,c,d)
Full E2E 5.19 12.63 10.98 21.27
Full Prompt 8.55 46.88 15.52 38.38
FlanT5-base 6.01 73.08 6.07 53.92
PEFT E2E 7.53 70.34 10.98 51.30
PEFT Prompt 9.11(a,b,c) 71.34(a,b,c,d) 12.98(a,b,c,d) 57.63(a,b,c,d)
Full E2E 7.42 44.76 10.91 31.22
Full Prompt 10.06 51.80 17.36 54.07
FlanT5-large 8.44 68.56 16.94 52.64
PEFT E2E 8.69 67.47 15.33 53.25
PEFT Prompt 11.24(a,b,d) 71.27(a,b,c,d) 18.17(a,b,c) 60.08(a,b,c,d)
Full E2E 11.75 27.31 14.79 23.17
Full Prompt 13.60 68.18 18.22 57.37
StableLM2 17.01 76.14 17.92 60.72
PEFT E2E 16.39 74.82 23.62 58.34
PEFT Prompt 21.75(a,b,c,d) 82.09(a,b,c,d) 27.23(a,b,c,d) 68.82(a,b,c,d)
Full E2E 14.23 5.87 13.22 2.46
Full Prompt 25.73 32.50 23.22 30.07
Table 3. Accuracy of vanilla and fine-tuned LMs, both with and without RAG. The RAG results are based on Ideal retrieval. The PEFT method is used for fine-tuning, and prompting is utilized for QA generation. The best results in each column are shown in bold, and the best results in each row are underlined. Statistically significant differences in the +FT+RAG columns are compared with the -FT+RAG columns. Superscript (a) indicates statistically significant differences (better or worse) as determined by the Wilcoxon test (p-value ¡ $0.01$).
Dataset PopQA WiTQA EQ
Model #P -FT +FT -FT +FT -FT +FT -FT +FT -FT +FT -FT +FT
-RAG -RAG +RAG +RAG -RAG -RAG +RAG +RAG -RAG -RAG +RAG +RAG
FlanT5-small 80m 2.69 7.26 47.46 61.39(a) 8.76 18.30 40.76 59.47(a) 2.84 9.39 27.36 42.68(a)
FlanT5-base 250m 6.01 9.11 73.08 71.34(a) 16.52 23.32 73.35 74.34 6.07 12.98 53.92 57.63(a)
FlanT5-large 780m 8.44 11.24 68.56 71.27(a) 24.52 28.85 74.37 77.24(a) 16.94 18.17 52.64 60.08(a)
Tiny-llama 1.1B 17.50 18.32 74.39 74.87 45.12 47.65 78.84 80.60(a) 21.12 24.57 61.03 61.61
StableLM2 1.6B 17.01 21.75 76.14 82.09(a) 42.18 51.66 81.08 86.19(a) 17.92 27.23 60.72 68.82(a)
MiniCPM 2B 14.16 15.47 69.44 75.86(a) 37.61 44.94 73.17 80.45(a) 15.31 22.92 54.63 62.67(a)
FlanT5-xl 3B 12.24 13.25 73.31 74.71(a) 31.24 36.38 76.97 79.67(a) 15.98 20.42 59.07 62.70(a)
Mistral 7B 21.47 30.70 80.25 78.44(a) 51.36 58.29 86.05 83.90(a) 25.78 34.01 68.60 64.96(a)
Zephyr 7B 28.23 35.48 78.65 78.20 58.33 63.74 83.83 86.89(a) 29.09 38.52 62.45 67.89(a)
Llama2-chat 7B 26.09 27.71 81.15 80.15 53.88 56.38 86.50 84.71(a) 27.13 32.84 68.29 67.11(a)
Llama3-chat 8B 32.52 32.75 81.29 81.54 61.88 61.58 86.86 85.71 35.07 37.95 68.67 68.64
FlanT5-xxl 11.3B 11.26 15.94 75.19 74.98 30.40 42.89 78.19 81.44(a) 12.48 23.16 59.67 62.80(a)
Datasets. We conduct our experiments on three datasets focused on factual knowledge: PopQA (Mallen et al., 2023), WitQA (Maekawa et al., 2024), and EntityQuestion (EQ) (Sciavolino et al., 2021), all of which include long-tail entities; see Table 1 for statistics. PopQA is an open-domain QA dataset about long-tail entiteis, constructed from 16 diverse relationship types in Wikidata. EQ is another popular open-domain QA dataset that covers a long-tail entity distribution, using Wikipedia hyperlink counts as a proxy for entity frequency and sampling knowledge triples from Wikidata based on these frequency distributions. Since EQ does not provide Wikidata IDs for each entity, we use only about 80% of the questions, where the mention of the subject entity has a unique match with a Wikidata entity. WitQA is another entity-centric dataset that defines a different proxy for popularity. They argue that the popularity metric should be based on the occurrence of both the subject entity and the relation (unlike the subject-based popularity in PopQA and EQ). Therefore, they define the S-R count, which is the co-occurrence of the subject entity and relation predicate. However, they report pageviews in their dataset, and we use the pageview-based popularity to enable comparison with other datasets. To analyze performance with respect to popularity, we divide the entities into five buckets based on their popularity levels; see Figure 5.
Evaluation Metric. Following previous studies (Mallen et al., 2023; Maekawa et al., 2024; Sciavolino et al., 2021), we report on the accuracy metric, where a prediction is considered correct if one of the ground truth responses matches a substring of the predicted response. While widely used (Kandpal et al., 2023; Liu et al., 2023; Cuconasu et al., 2024), this metric is not without issues. A principal problem arises in determining response correctness, particularly in cases involving date representations or varying phrasings that convey identical meanings (Cuconasu et al., 2024). For example, comparing the predicted response “Nathanson” with ground truth “Jeff Nathanson,” the prediction is considered incorrect. Another observed problem is that when the model generates multiple entity names, among them the ground truth entity, the response is incorrectly considered as correct. Recognizing these limitations, we acknowledge the necessity for a more advanced analysis of answer variations, which we leave for future research.
Table 4. Performance of retrievers. The relevant document for each question is assumed to be the first paragraph of the corresponding Wikipedia page. Statistically significant differences in the DPR row are compared with other rows. Superscripts (a), (b), and (c) indicate statistically significant differences (better or worse) compared to BM25, Contriever, and BM25+DPR, respectively, as determined by a t-test (p-value ¡ $0.01$).
Dataset PopQA WiTQA EQ
Model Rec@1 Rec@3 Rec@5 Rec@1 Rec@3 Rec@5 Rec@1 Rec@3 Rec@5
BM25 40.13 62.92 71.30 56.40 81.88 88.87 64.01 89.00 93.71
Contriever 42.90 70.84 80.17 53.23 84.25 91.52 63.10 91.93 96.30
BM25+DPR 59.35 80.85 87.55 72.01 90.17 93.72 79.10 95.09 96.11
DPR 59.40(a, b) 80.85(a, b) 87.57(a, b) 71.73(a, b) 90.05(a, b) 93.62(a, b) 79.01(a, b) 94.06(a, b) 96.10(a)
Table 5. Overall accuracy of answer generation using different retrievers before and after FT. The performance of LMs in answering questions correlates with the effectiveness of the retrieval models. Statistically significant differences in the DPR rows are compared with other rows. Superscripts (a), (b), and (c) indicate statistically significant differences (better or worse) compared to BM25, Contriever, and BM25+DPR, respectively, as determined by the Wilcoxon test (p-value ¡ $0.01$).
PopQA WiTQA EQ
Model BM25 Contr- iever BM25+ DPR DPR BM25 Contr- iever BM25+ DPR DPR BM25 Contr- iever BM25+ DPR DPR
FlanT5-base -FT 41.40 44.48 54.21 53.36(a, b) 60.25 58.06 66.62 66.51(a, b) 43.90 43.01 49.62 49.54(a, b)
FlanT5-base +FT 39.99 43.24 52.19 52.21(a, b) 61.06 59.05 66.89 66.83(a, b) 45.35 45.24 50.45 50.39(a, b)
StableLM2 -FT 46.05 47.83 55.49 55.74(a, b) 72.52 69.33 74.53 74.47(a, b) 51.71 49.25 54.60 54.51(a, b)
StableLM2 +FT 49.92 53.63 60.97 61.44(a, b) 77.67 76.65 80.48 80.45(a, b) 59.99 59.24 63.67 61.91(a, b)
Mistral -FT 49.58 53.10 60.14 60.09(a, b) 78.33 77.35 81.54 81.49(a, b) 59.09 58.03 63.13 63.06(a, b)
Mistral +FT 50.53 54.08 61.38 61.23(a, b) 76.82 75.60 79.35 79.22(a, b) 56.74 55.19 61.43 61.28(a, b)
Llama3 -FT 51.30 56.49 61.57 61.46(a, b) 78.51 79.00 81.62 81.73(a, b) 59.31 59.26 63.37 63.27(a, b)
Llama3 +FT 53.57 56.90 62.05 62.10(a, b) 79.96 79.59 81.72 81.77(a, b) 61.86 61.05 64.07 64.00(a, b)
Language Models. We use several LMs, focusing on two main features: the backbone architecture (i.e., decoder-only and encoder-decoder) and model size, which ranges from 80 million to over 11 billion parameters. For the encoder-decoder models, we utilize five versions of the FlanT5 family (Chung et al., 2022), spanning from small to XXL. For the decoder-only models, we employ smaller LMs such as Tiny-llama (Zhang et al., 2024b), StableLM2 (Bellagente et al., 2024), and MiniCPM (Hu et al., 2024), which range from 1 to 2 billion parameters. Additionally, we incorporate larger LMs like Mistral (Jiang et al., 2023), Zephyr (Tunstall et al., 2023), LLama2 (Touvron et al., 2023), and LLama3, [^1] https://ai.meta.com/blog/meta-llama-3 which range from 7 to 8 billion parameters. As for instructions, we apply zero-shot prompting for generative prediction using a straightforward templates. For non retrieved-augmented input, we use template: "Question: <question>", and for retrieved-augmented input, we use template: "Context: <context> Question: <question>".
Fine Tuning. We generate training data for the FT approach using two distinct data augmentation methods. To ensure a fair comparison between FT and RAG with an ideal retriever, we generate QAs exclusively using the summary sections of Wikipedia pages. After generating QA pairs, we proceed to fine-tune LMs using two approaches: full parameter tuning (Full) and Parameter Efficient Fine-Tuning (PEFT). Within the range of PEFT techniques (Zaken et al., 2022; Liu et al., 2022; Ma et al., 2023), we utilize QLoRA (Dettmers et al., 2023), chosen for its broad acceptance in the field (Kaddour et al., 2023; Naveed et al., 2023) and efficient use of computational resources we had at our disposal.
RAG. We utilize a variety of retrieval models to obtain relevant documents for the RAG approach, including BM25 (Robertson and Zaragoza, 2009), Contriever (Izacard et al., 2022), DPR (Karpukhin et al., 2020), and a two-stage re-ranker that combines BM25 with DPR, all implemented according to the BEIR benchmark (Thakur et al., 2021). Additionally, since the selected datasets do not contain grounded document evidence, we assume that the summary section of each Wikipedia page is the answer-containing document. We define an ideal retriever model as one that returns the summary paragraph as the top-ranked document, referred to as the Ideal retriever throughout the paper. We acknowledge that this assumption is not entirely accurate as some answers may be found in other subsections. However, our evaluation on the downstream task (Figures 6 and 8) demonstrates that the Ideal retriever outperforms other retrievers.
Stimulus RAG. We select a DPR retriever for the hint extractor and set $K=3$. We report on two variation of the SRAG approach: (i) SRAG(S), which utilizes the top-1 sentence as the hint, and (ii) SRAG(D), which inserts the entire document contacting the top-ranked sentence. For the instruction, we used the following template: "Context: <hint><context> Question: <question>".
5. Results
In the following, we evaluate fine-tuning and RAG methods on different setups and answer our two research questions listed in Section 1.
5.1. Fine Tuning vs. RAG Performance
The first research question (RQ1) involves the comparison between FT and RAG methods and factors affecting these models: (i) fine-tuning method, (ii) data augmentation method, (iii) LM type and size, and (iv) retrieval model performance.
Comparison of fine tuning methods. To study the effect of fine-tuning method, we investigate four LMs: three from the FlanT5 family (encoder-decoder models) and StableLM2 (a decoder-only model). These models are chosen as their parameters are less than 2B, allowing us to perform full fine-tuning with resources we had at our disposal. Table 2 shows that PEFT leads to smaller gains in the +FT-RAG setup compared to full FT in most cases, yet it significantly improves accuracy in the +FT+RAG setup. This suggests that PEFT enables the LM to maintain its reasoning abilities based on the provided prompts. Based on this observation, we selected PEFT as the fine-tuning method for subsequent experiments.
Comparison of data augmentation methods. The E2E method generates over 12 times more QAs than the prompting method, while the prompt-based method generates higher quality QAs. The results in Table 2 show that fine-tuned models trained on prompt-generated data outperform those trained on E2E-generated data. This highlights the significance of synthetic data quality over quantity. As a result, for subsequent experiments, we use QAs generated by the prompt method.
Comparison of FT and RAG with LMs of different type and size. Choosing PEFT as the FT method and the Prompt method for QA generation, we extended our experiments to 12 LMs using the ideal retriever for RAG. Table 3 presents the results. The findings indicate that while FT enhances the answer generator’s performance, it alone cannot match or approach the performance of RAG. However, combining FT with RAG yields the best results for smaller models (up to 3B parameters). Conversely, for larger models (from 7B to 11.3B parameters), combining FT with RAG degrades performance. This suggests that although FT injects knowledge into LMs (as seen when comparing -FT-RAG with +FT-RAG), it diminishes reasoning abilities of larger LMs. Interestingly, since smaller LMs’ accuracy improves with FT, the best result for +FT+RAG is achieved by StableLM2, which has only 1.6B parameters. Another observation is that decoder-only models perform better than encoder-decoder (Flan-T5) models of similar size.
Figure 6. The performance of the LMs with different combinations of FT and RAG. The +FT+RAG setup outperforms other setups across all models and datasets.
Effect of retrieval models on RAG. To assess the effect of retrieval models, we first calculate the overall Recall score for all retriever models, considering the first paragraph of each Wikipedia page as the answer-containing document. The results are shown in Table 4, demonstrating that DPR and BM25+DPR models significantly outperform BM25 and Contriever across all datasets, while DPR and BM25+DPR perform on par with other. Table 5 presents the QA accuracy of RAG with different retrieval models. Following Mallen et al. (2023), we use the top-ranked document for the RAG prompt. The results indicate a direct correlation between the performance of the retrieval model and the overall QA accuracy, underscoring the significant impact of the retrieval model on the effectiveness of the downstream task for both vanilla and fine-tuned LMs. As performance of RAG with DPR and BM25+DBPR is comparable, we use DPR as the retrieval model for subsequent experiments.
Analysis of RAG and FT per popularity. Figure 6 illustrates QA accuracy across different popularity buckets, providing insight into the effectiveness of FT and RAG (using Ideal retriever). It is evident that RAG significantly increases accuracy for the least popular entities, which aligns with the findings of Mallen et al. (2023). Additionally, these figures demonstrate that FT enhances QA accuracy across all popularity levels, with the most notable improvements observed in the least popular buckets for Mistral and Llama3.
Analysis of retrieval models per popularity. Figure 7 compares the performance of retrieval models against the Ideal retriever across different popularity buckets. The results indicate that retrieval effectiveness is higher for less popular entities compared to more popular entities. This is likely due to the limited occurrences of noisy documents for less popular entities.
Figure 8 shows the QA system’s accuracy using various retrieval models within the RAG framework across different popularity buckets for the FlanT5-base and StableLM2 models. The left figures display results for vanilla LMs, while the right figures show results for fine-tuned LMs. FT does not alter the pattern of accuracy across popularity buckets but shifts the overall accuracy higher. Interestingly, the accuracy decreases from the less popular bucket to the fourth popularity bucket across different retrievers but increases in the most popular bucket. This reduction in accuracy can be interpreted by the finding in Figure 7, which shows that the retriever’s performance decreases as popularity increases. However, it appears that for popular entities, the LMs can ignore noisy information in the input prompt and rely on their internal knowledge to answer questions, resulting in a sudden increase in accuracy despite the retriever’s lower performance in the most popular bucket.
Figure 7. Recall@1 for retrieval models across different popularity levels shows that retrievers perform more effectively with less popular knowledge compared to more popular ones.
Figure 8. Performance of the answer generator task across popularity buckets on PopQA. FT does not alter the overall pattern. Accuracy decreases as the retriever models’ performance drops from the least popular bucket to the fourth bucket. Interestingly, accuracy increases for the most popular bucket, indicating that LMs rely on their embedded information for popular entities.
Table 6. The effect of increasing the number of documents in RAG. A significant jump in accuracy is observed when the number of documents is increased to three. However, adding five documents does not significantly affect accuracy. Statistically significant differences in the (3D) columns are compared with the (1D) and (5D) columns. Superscripts (a) and (b) indicate statistically significant differences (better or worse) compared to (1D) and (5D), respectively, as determined by the Wilcoxon test (p-value ¡ $0.01$).
-FT+RAG +FT+RAG
Model (1D) (3D) (5D) (1D) (3D) (5D)
PopQA
FlanT5-base DPR 53.36 56.67(a) 56.67 52.20 53.46(a) 53.46
Ideal 74.50 73.06(b) 75.53 71.34 72.02 71.75
StableLM2 DPR 55.74 63.98(a) 64.00 61.44 65.33(a) 65.33
Ideal 76.14 80.82(a) 80.68 82.09 82.98 82.99
Mistral DPR 60.09 65.22(a) 65.22 61.23 63.63(a) 63.63
Ideal 80.25 81.58(a) 81.43 78.44 80.30(a) 80.44
Llama3 DPR 61.46 66.66(a,b) 67.94 62.10 66.61(a) 66.43
Ideal 81.29 82.58(a) 83.27 81.54 82.58 82.63
EQ
FlanT5-base DPR 49.54 52.39(a) 52.05 50.39 48.90(a) 48.02
Ideal 53.92 58.02(a) 58.18 57.63 57.24 57.52
StableLM2 DPR 54.51 63.51(a) 63.75 61.91 64.67(a) 64.52
Ideal 60.72 69.16(a) 69.16 68.82 70.74(a) 71.09
Mistral DPR 54.51 66.85(a) 67.40 61.28 63.92(a) 64.73
Ideal 68.60 71.38(a) 71.98 64.94 68.60(a,b) 70.31
Llama3 DPR 63.27 67.76(a) 68.41 64.00 65.56(a) 65.45
Ideal 68.67 71.50(a) 72.46 68.64 71.89(a) 71.81
5.2. Stimulus RAG performance
The second research question (RQ2) concerns whether our proposed Stimulus RAG method can surpass the performance of fine-tuned models. To evaluate the effect of the stimulus RAG, we first need to investigate how increasing the number of documents in the input prompt affects the accuracy of LMs. Table 6 presents the results of RAG with top-1, top-3 and top-5 documents, shown as (1D), (3D), and (5D), respectively. It shows that using top-3 documents leads to noticeable accuracy improvements in all cases, both before and after FT. For DPR, these results align with those in Table 4, where there is a significant increase from Recall@1 to Recall@3. For Ideal retriever, it is important to note that the Ideal retriever is not 100 percent accurate; for some queries, the answer is found in other paragraphs, not just in the summary paragraph.
Another notable observation is that increasing the number of input documents to five results in either negligible accuracy improvement or a decrease in accuracy. This occurs despite Table 7 showing that Recall@5 for DPR is higher than Recall@3. This observation suggests that as the number of documents increases, some LMs struggle to effectively utilize all the information, making it harder to find the correct answer and leading to misunderstandings. The results indicate that adding more textual information to the input prompt should be done judiciously.
Table 7. SRAG performance. By adding the extracted hint to the top of the input prompt, SRAG’s performance surpasses other settings. Statistically significant differences in the SRAG(S) and SRAG(D) columns are compared with the -FT+RAG and +FT+RAG columns. Superscripts (a) and (b) denote statistically significant differences (better or worse) compared to -FT+RAG and +FT+RAG, respectively, as determined by the Wilcoxon test (p-value ¡ $0.01$).
-FT+RAG +FT+RAG SRAG
Model (3D) (3D) (S) (D)
PopQA
FlanT5-base DPR 56.67 53.46 57.77(a,b) 57.67(a,b)
Ideal 73.06 72.02 75.08(a,b) 75.29(a,b)
StableLM2 DPR 63.98 65.33 65.48(a) 66.01(a)
Ideal 80.82 82.98 82.83(a) 83.18(a)
Mistral DPR 65.22 63.63 65.84(a,b) 66.04(a,b)
Ideal 81.58 80.30 81.88(b) 82.27(a,b)
Llama3 DPR 66.66 66.61 67.22 67.21
Ideal 82.58 82.58 82.42 81.60
EQ
FlanT5-base DPR 52.39 48.90 55.31(a,b) 55.12(a,b)
Ideal 58.02 57.24 60.21(a,b) 59.88(a,b)
StableLM2 DPR 63.51 64.67 65.97(a,b) 66.73(a,b)
Ideal 69.16 70.74 71.38(a) 71.73(a,b)
Mistral DPR 66.85 63.92 68.54(a,b) 68.76(a,b)
Ideal 71.38 68.60 72.23(a,b) 72.45(a,b)
Llama3 DPR 67.76 65.56 68.31(b) 68.38(b)
Ideal 71.70 71.89 71.71 71.98
Table 7 presents the results of the Stimulus RAG with three documents. It is important to note that the hint sentence and document are derived from the top-3 documents, so no extra information is added to the input prompt. This ensures a fair comparison with RAG using three documents, allowing us to solely evaluate the effectiveness of the highlighting method. In all cases, Stimulus RAG without FT achieves higher accuracy than fine-tuned LMs with RAG using top-3 documents. This indicates that guiding LMs with a hint not only improves RAG accuracy before FT but can also surpass the effects of FT. These findings demonstrate that better accuracy can be achieved by designing a more advanced RAG system without the complexities and resource demands of FT.
6. Discussion and Conclusions
In this paper, we aimed to determine the most suitable approach for customizing language models (LMs) for less-resourced domains. We examined the effectiveness of retrieval augmented generation (RAG) and fine-tuning (FT) methods, focusing on four key aspects: (i) fine-tuning methods, specifically full fine-tuning versus parameter-efficient fine-tuning (PEFT), (ii) data augmentation techniques, (iii) the type and size of LMs, including decoder-only versus encoder-decoder models ranging from 80 million to 11 billion parameters, and (iv) the performance of retrieval models. Our findings reveal several key points. First, PEFT enhances downstream task performance and preserves the reasoning abilities of LMs while incorporating new knowledge. Second, prompt-based QA generation exhibits superior performance in factual QA tasks. Third, a small fine-tuned LM with RAG can perform on par with or even surpass a larger LM model. Additionally, RAG’s performance improves with higher-performing retrievers. Notably, when comparing knowledge injection methods, RAG significantly outperforms FT. We addressed the cost of fine-tuning by developing Stimulus RAG (SRAG), a novel RAG approach that prompts an LM to generate correct responses based on hints provided in the prompt. This method eliminates the need for extensive fine-tuning, making it a cost-effective solution for enhancing LM performance in less-resourced domains.
Acknowledgments
This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).